# Don’t Waste It: Guiding Generative Recommenders with Structured Human Priors via Multi-head Decoding

### v1

@@ 1 Introduction
The goal of recommender systems extends beyond mere predictive accuracy. The importance of objectives such as novelty and diversity has long been recognized within the academic community adomavicius2005toward, acknowledging that a successful system must balance relevance with discovery. Nevertheless, the metrics predominantly used to evaluate and optimize these systems have centered on accuracy and engagement ricci2021recommender. This focus catalyzed significant algorithmic advancements, from collaborative filtering sarwar2001item; he2017neural to deep sequential models kang2018self; zhai_actions_2024. However, the prioritization of easily measurable signals has exposed a critical limitation, often termed the alignment problem: a model may predict the next interaction accurately, yet fail to align with the user’s broader goals or well-being. This optimization imbalance has been shown to yield detrimental side effects, including polarization, addiction, and popularity bias, while discouraging the discovery of new user interests. This realization has accelerated a paradigm shift toward a human-centered approach. The critical question is evolving from “Is this recommendation accurate?” to “Is this recommendation worth of your time?”, which requires considering a richer set of objectives that extend beyond accuracy, such as diversity, novelty, and personalization said2025weredoingallwrong.
@@
This dichotomy between complex post-hoc adjustments and unsupervised E2E models motivates a question: Instead of discarding the human priors accumulated over years of practice, can we integrate them directly into the learning process of generative recommender systems in a simple, effective, and interpretable manner? To this end, we propose a backbone-agnostic framework that seamlessly injects various human prior into the generative model training with lightweight adapters, by drawing inspiration from efficient decoding strategies in Large Language Models (LLMs) cai2024medusasimplellminference. Unlike post-hoc filtering or architecture-specific modifications, these adapter heads guide the sequential model to learn user representations that are naturally disentangled. This renders the model inherently controllable, explainable, and better aligned with complex, real-world objectives.
@@ 3.2 Incorporating Human Priors via Conditioned Query Heads
With compatibility masking, different query heads are specialized with distinct functional roles determined by the specified priors. Thus, when predicting for one prior group, our method can largely reduce the reliance on irrelevant features, which minimizes the mutual interference among these different objectives. As a result, the model can show stronger predictive capability for each prior group, and thus result in an performance improvement with the cooperation of the heads.
@@ 4.3 Main Results
Table 2: Overall performance comparison. Human priors consistently lead to improvements over the backbone models (HSTU and HLLM). The backbones and baselines are highlighted in gray. Note that HLLM and HSTU results are not directly comparable due to different context lengths used (See Appendix 6.1), and all the baselines are ran under the HSTU settings and should only be compared to HSTU.
@@ 4.4.2 User Interest Exploration
Exploration in recommender systems aims to uncover content a user may like but has not yet engaged with. while it improves long-term user engagement, it is often believed to hurt near-term experience chenexplorationrs2021, similar to the exploration versus exploitation dilemma in reinforcement learning. We argue that adding multiple heads for human priors can somehow strike a good balance between the two. To evaluate this, we analyze the performance of our method on a targeted subset of users in the Pixel8M dataset using HLLM. Specifically, we identify users who interacted with item features during the forecast horizon of the test split, but who had never engaged with those features in their prior history. In these cases, the model’s success hinges on its ability to recommend items from entirely new categories. Applying this criterion, we identified a total of 283,497 such users.
@@ 4.4.3 Personalization
Another common issue in recommender systems is popularity bias. In some cases, it means that the behaviors of some users, whose interests are different from the majority, might get inferior recommendations. To address this issue, we instantiated User Prior on EB-NeRD, constructed similar to Graph Prior. But here, the nodes are the users in the co-engagement graph, and an edge exists between the two users if they interacted with the same item in the train set. To avoid popular items from making the graph too dense, for each item, we random sample a maximum of 2000 users that have interacted with it before generating the edges. We then employ the Leiden algorithm Traag2018FromLT with the modularity objective to cluster the graph into a total of 9 user groups. We define User Prior by assigning an adapter to each group.
@@
As shown in Figure 2, before introducing any prior, the minority groups with fewer users generally suffer from lower recommendation quality. However, by dedicating a separate head to each user group, we effectively lower the impact of the majority users from dominating the query representation, which allows us to better personalize the recommendations for the minority users. This results in bigger improvements for users with fewer users, and recommendation quality looks more balanced after User Prior is introduced.
@@ 4.5 Scalability
A key trend in recommender systems is the use of longer context lengths to enhance personalization. While scaling laws typically require more training data to improve larger models, our experimental setup with a fixed dataset introduces a trade-off: increasing the context length reduces the number of available training windows. To investigate this, we test HSTU with longer context lengths and larger model sizes on Pixel8M. In Figure 3, the left subplot shows that for size 4 and 1b, the base model (LT/ST1, equivalent to the discounted loss) struggles to benefit from increasing context as we increase the context length beyond 20 items. However, in the right subplot, when guided by human priors, the same model architecture continues to extract performance gains from longer contexts and larger model sizes, although the magnitude of increase slowly plateaus. This finding suggests that the structural information imposed by human priors facilitates more efficient learning, allowing the model better benefit from increasing context and larger model sizes when the amount of training data is fixed.
@@ 4.6 Representation Space Visualization
We observe that the recommendations of “No Prior” form a single, concentrated cluster, failing to span the user’s diverse targets. In contrast, ”w/ Item Prior” leverages its different heads to generate recommendations that cover the full spectrum of the target items. For REMI, we set the number of interest to two since it achieves the best performance. However, one interest (orange squares) dominates the recommendations in the bottom-left corner with fewer target items, while the second interest (light blue squares) only partially covers the primary target cluster in the upper-right corner. This suggests that leveraging existing human prior achieves a more effective alignment between its diverse heads and the user’s multifaceted interests compared to learning the multiple interests as in REMI.
@@ 4.7.2 Why it works: priors or just more heads?
Table 5: Average gain over the baseline HLLM (w/o prior) across eight metrics. We train the model train on $\tau=4$, and show its robustness when evaluated on $\tau=1,4,8$.
@@ 5 Conclusion
By integrating human priors accumulated over post-hoc business rules directly into generative recommender’s learning process, our proposed framework offers a principled approach to aligning recommendations with multifaceted human objectives. Our experiments demonstrate that prior-conditioned adapter heads not only enhance accuracy, but also improve diversity, novelty, and personalization, which are key dimensions of user experience that are often overlooked. The framework’s backbone-agnostic design and hierarchical composition strategy further enable flexible modeling of complex user intent, making it broadly applicable across scenarios.
@@ 6.2 Model Architectures and Hyperparameters
For both HSTU and HLLM backbones, we use a discount factor $\gamma=0.99$. However, we observed that a smaller $\gamma$ might lead to higher accuracy at the cost of lower diversity (Table 7). r the green items in the first three target items.
@@
Baseline Configurations. For ComiRec and REMI, we use the self-attention version, which demonstrates performance comparable to the dynamic routing version, but with a stabler and faster training. We use a learning rate of $0.001$, search the number of interest from $\{1,2,4,8,16,32\}$, and report the best performance. For REMI, we tune the additional $\beta$ parameter for interest-aware hard negative mining from $\{0.1,1,4,10\}$ and set routing regularization weight $\lambda$ to $100$ following the original paper.

### v2

@@ 1 Introduction
The goal of recommender systems extends beyond mere predictive accuracy. The importance of objectives such as novelty and diversity has long been recognized within the academic community adomavicius2005toward, acknowledging that a successful system must balance relevance with discovery. Nevertheless, the metrics predominantly used to evaluate and optimize these systems have centered on accuracy and engagement ricci2021recommender. This focus catalyzed significant algorithmic advancements, from collaborative filtering sarwar2001item; he2017neural to deep sequential models kang2018self; zhai_actions_2024. However, the prioritization of easily measurable signals has exposed a critical limitation, often termed the alignment problem: a model may predict the next interaction accurately, yet fail to align with the user’s broader goals or well-being. This optimization imbalance has been shown to yield detrimental side effects, including polarization, addiction, and popularity bias, while discouraging the discovery of new user interests. This realization has accelerated a paradigm shift toward a human-centered approach. The critical question is evolving from “Is this recommendation accurate?” to “Is this recommendation worth your time?”, which requires considering a richer set of objectives that extend beyond accuracy, such as diversity, novelty, and personalization said2025weredoingallwrong.
@@
This dichotomy between complex post-hoc adjustments and unsupervised E2E models motivates a question: Instead of discarding the human priors accumulated over years of practice, can we integrate them directly into the learning process of generative recommender systems in a simple, effective, and interpretable manner? To this end, we propose a backbone-agnostic framework that seamlessly injects various human priors into the generative model training with lightweight adapters, by drawing inspiration from efficient decoding strategies in Large Language Models (LLMs) cai2024medusasimplellminference. Unlike post-hoc filtering or architecture-specific modifications, these adapter heads guide the sequential model to learn user representations that are naturally disentangled. This renders the model inherently controllable, explainable, and better aligned with complex, real-world objectives.
@@ 3.2 Incorporating Human Priors via Conditioned Query Heads
With compatibility masking, different query heads are specialized with distinct functional roles determined by the specified priors. Thus, when predicting for one prior group, our method can largely reduce the reliance on irrelevant features, which minimizes the mutual interference among these different objectives. As a result, the model can show stronger predictive capability for each prior group, and thus result in a performance improvement with the cooperation of the heads.
@@ 4.3 Main Results
Table 2: Overall performance comparison. Human priors consistently lead to improvements over the backbone models (HSTU and HLLM). The backbones and baselines are highlighted in gray. Note that HLLM and HSTU results are not directly comparable due to different context lengths used (See Appendix 7.1), and all the baselines are run under the HSTU settings and should only be compared to HSTU.
@@ 4.4.2 User Interest Exploration
Exploration in recommender systems aims to uncover content a user may like but has not yet engaged with. Although it improves long-term user engagement, it is often believed to negatively affect near-term experience chenexplorationrs2021, similar to the exploration versus exploitation dilemma in reinforcement learning. We argue that adding multiple heads for human priors can somehow strike a good balance between the two. To evaluate this, we analyze the performance of our method on a targeted subset of users in the Pixel8M dataset using HLLM. Specifically, we identify users who interacted with item features during the forecast horizon of the test split, but who had never engaged with those features in their prior history. In these cases, the model’s success hinges on its ability to recommend items from entirely new categories. Applying this criterion, we identified a total of 283,497 such users.
@@ 4.4.3 Personalization
Another common issue in recommender systems is popularity bias. In some cases, it means that the behaviors of some users, whose interests are different from the majority, might receive inferior recommendations. To address this issue, we instantiated User Prior on EB-NeRD, constructed similar to Graph Prior. But here, the nodes are the users in the co-engagement graph, and an edge exists between the two users if they interacted with the same item in the train set. To avoid popular items from making the graph too dense, for each item, we random sample a maximum of 2000 users that have interacted with it before generating the edges. We then employ the Leiden algorithm Traag2018FromLT with the modularity objective to cluster the graph into a total of 9 user groups. We define User Prior by assigning an adapter to each group.
@@
As shown in Figure 2, before introducing any prior, the minority groups with fewer users generally suffer from lower recommendation quality. However, by dedicating a separate head to each user group, we effectively lower the impact of the majority users from dominating the query representation, which allows us to better personalize the recommendations for the minority users. This results in bigger improvements for groups with fewer users, and recommendation quality looks more balanced after User Prior is introduced.
@@ 4.5 Scalability
A key trend in recommender systems is the use of longer context lengths to enhance personalization. While scaling laws typically require more training data to improve larger models, our experimental setup with a fixed dataset introduces a trade-off: increasing the context length reduces the number of available training windows. To investigate this, we test HSTU with longer context lengths and larger model sizes on Pixel8M. In Figure 3, the left subplot shows that for size 4 and 1b, the base model (LT/ST1, equivalent to the discounted loss) struggles to benefit from increasing context as we increase the context length beyond 20 items. However, in the right subplot, when guided by human priors, the same model architecture continues to extract performance gains from longer contexts and larger model sizes, although the magnitude of increase slowly plateaus. This finding suggests that the structural information imposed by human priors facilitates more efficient learning, allowing the model to better benefit from increasing context and larger model sizes when the amount of training data is fixed.
@@ 4.6 Representation Space Visualization
We observe that the recommendations of “No Prior” form a single, concentrated cluster, failing to span the user’s diverse targets. In contrast, ”w/ Item Prior” leverages its different heads to generate recommendations that cover the full spectrum of the target items. For REMI, we set the number of interest to two since it achieves the best performance. However, one interest (orange squares) dominates the recommendations in the bottom-left corner with fewer target items, while the second interest (light blue squares) only partially covers the primary target cluster in the upper-right corner. This suggests that leveraging existing human priors achieves a more effective alignment between its diverse heads and the user’s multifaceted interests compared to learning the multiple interests as in REMI.
@@ 4.7.2 Why it works: priors or just more heads?
Table 5: Average gain over the baseline HLLM (w/o prior) across eight metrics. We train the model on $\tau=4$, and show its robustness when evaluated on $\tau=1,4,8$.
@@ 5 Conclusion
By integrating human priors accumulated over post-hoc business rules directly into generative recommenders’ learning process, our proposed framework offers a principled approach to aligning recommendations with multifaceted human objectives. Our experiments demonstrate that prior-conditioned adapter heads not only enhance accuracy, but also improve diversity, novelty, and personalization, which are key dimensions of user experience that are often overlooked. The framework’s backbone-agnostic design and hierarchical composition strategy further enable flexible modeling of complex user intent, making it broadly applicable across scenarios.
@@
6 Acknowledgment
We would like to thank Zhuokai Zhao, Benyu Zhang, Lizhu Zhang, Xiangjun Fan, and Abhishek Kumar for their support.
@@ 7.2 Model Architectures and Hyperparameters
For both HSTU and HLLM backbones, we use a discount factor $\gamma=0.99$. However, we observed that a smaller $\gamma$ might lead to higher accuracy at the cost of lower diversity (Table 7).
@@
Baseline Configurations. For ComiRec and REMI, we use the self-attention version, which demonstrates performance comparable to the dynamic routing version, but with stabler and faster training. We use a learning rate of $0.001$, search the number of interest from $\{1,2,4,8,16,32\}$, and report the best performance. For REMI, we tune the additional $\beta$ parameter for interest-aware hard negative mining from $\{0.1,1,4,10\}$ and set routing regularization weight $\lambda$ to $100$ following the original paper.